# The quantum metric of electrons with spin-momentum locking

### v1

@@ Acknowledgments
We thank Jean-Marc Triscone, Marc Gabay, Ajit Srivastava, Edouard Lesne, and Graham Kimbell for fruitful discussions. We thank Marco Lopes and Jean-Luc Lorenzoni for technical support. We thank Alberto Morpurgo and Ignacio Gutiérrez for contributing to the sample fabrication. Funding G.S. and A.D.C. acknowledge ERC/SERI, Moore Foundation. M.T.M and M.C. acknowledge support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE). M.T.M. and C.O. acknowledge support from the MAECI project ”ULTRAQMAT”. M.C. acknowledges financial support from PNRR MUR project PE0000023-NQSTI. Author contributions G.S. fabricated the oxide heterostructures and the transport devices used in this study. G.S. and K.D. performed the measurements and analyzed the data with the support of S.G.. M.T.M., M.C., and C.O. developed the theoretical model. A.D.C. supervised the experimental work, and C.O. supervised the theoretical work. All authors contributed to writing the manuscript. Competing interests The authors declare no competing financial interests. Data and materials availability The data that support the findings of this study are available from Zenodo  (?) and from the corresponding authors upon reasonable request.

### v2

@@ Acknowledgments
We thank Jean-Marc Triscone, Marc Gabay, Ajit Srivastava, Edouard Lesne, and Graham Kimbell for fruitful discussions. We thank Marco Lopes and Jean-Luc Lorenzoni for technical support. We thank Alberto Morpurgo and Ignacio Gutiérrez for contributing to the sample fabrication. Funding This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract no. MB22.00071, by the Gordon and Betty Moore Foundation (grant no. 332 GBMF10451 to A.D.C.), by the European Research Council (ERC), and by the Dutch Research Council (NWO) as part of the VIDI (project 016.Vidi.189.061 to A.D.C.). M.T.M and M.C. acknowledge support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE). M.T.M. and C.O. acknowledge support from the MAECI project ”ULTRAQMAT”. M.C. acknowledges financial support from PNRR MUR project PE0000023-NQSTI. Author contributions G.S. fabricated the oxide heterostructures and the transport devices used in this study. G.S. and K.D. performed the measurements and analyzed the data with the support of S.G.. M.T.M., M.C., and C.O. developed the theoretical model. A.D.C. supervised the experimental work, and C.O. supervised the theoretical work. All authors contributed to writing the manuscript. Competing interests The authors declare no competing financial interests. Data and materials availability The data that support the findings of this study are available from Zenodo  (?) and from the corresponding authors upon reasonable request.